4. (5pts ) Write equation of the tangent line to the graph of f(x) = 5 - 2x at x = 0 Answer:
Answer #1

s f(x) = then Soun x = 0 at f(x) = 5- 22 so point (xo, yo) = (0,5) - 2 at x=0 slope (m) = f(x) m (x-xo) line is given by :y-